OpenStudy (agl202):

Wanna help?!? Get a MEDAL!!! What is the f of x over the g of x when f(x) = − 2x3 + x2 + 16x − 15 and g(x) = x + 3? −2x2 − 5x + 1 −2x2 + 7x − 5 −2x2 + 7x − 5 − 12 over the quantity of x plus 3 −2x2 − 5x + 1 − 12 over the quantity of x plus 3

OpenStudy (kpop4life123):

factor out x+2 from 7x2+14x and note that x3+8 = x3+2^3 = (x+2)(x2-2x+4)

OpenStudy (agl202):

Nvm.. I got it. It's B. Thx anyways! ;)